Illustrates the concepts pertinent to this problem. A refrigerator has a surfaces area of 5.3m^2. It is lined with 0.075-m-thick insulation whose thermal conductivity ia 0.030J/(s.m.C°). The interior temperature is kept at 5°C, while the temperature at the outside surface is 25°C. How much heat per second is being removed from the unit?
1
Expert's answer
2018-08-20T10:12:09-0400
dQ/dt=kA(T_2-T_1 )/d
The heat per second is
Q=dQ/dt ∆t=(0.03)(5.3)(25-5)/0.075 (1)=42 J.